OpenStudy (anonymous):

absolute value of 2x+8=-16 please solve

OpenStudy (lgbasallote):

simple trick to remember... |2x + 8| = -16 to get that you simply solve for x when 2x + 8 = -16 and 2x + 8 = -(-16) got it?

OpenStudy (anonymous):

Also, separate the x onto one side. (Subtract 8 from each side of the equations Igbasallote added, then divide by two) This will give you the value of each individual x.